The motor driving a grindstone is switched off when the latter has a rotational speed of 240 rpm. After 10sec, the speed is 180 rpm. If the angular retardation remains constant, how many additional revolutions does it make before coming to rest?


1
Expert's answer
2021-07-14T05:48:43-0400

"f_i= \frac{240}{60}=4 rps \\
f_i= \frac{180}{60}=3 rps \\
w_f=2\pi(3)=6\pi rad/s \\
w_f=w_i+\alpha t \\
6 \pi=2 \pi+\alpha *10 \implies \alpha=-0.2 \pi rad/s^2 \\
w_f^2=w_i^2+2\alpha \theta \\
0^2=(6\pi)^2+2(-0.2\pi) \theta \implies \theta =\frac{36\pi^2}{0.4\pi}=90\pi rad\\ 
N= \frac{90\pi}{2\pi}=45 rev"